OpenStudy (anonymous):

3.5/5 =.7 but what is that in a fraction

OpenStudy (anonymous):

.7 is in the tens place so it is 7/10

OpenStudy (anonymous):

\[ (7/2) = 3.5\] \[ \rightarrow \frac{3.5}{5} = \frac{7/2}{5} = \frac{7}{2*5} = \frac{7}{10} \]